# DTCLMapper: Dual Temporal Consistent Learning for Vectorized HD Map Construction

### v1

@@ III-B Temporal Instance Consistent Learning
As mentioned above, most previous works [8, 21, 4], use temporal information to improve the richness of visual features in BEV space. In their works, global BEV features, as the basic unit, are analyzed through convolution or attention mechanism to get the global weight matrix in temporal frames. Then, this matrix is leveraged to fuse the missing features of this frame from the historical BEV features. Since the environment structure in continuous frames is very consistent, together with this complementary fusion of differential features, the similar features distributed in a large range will also be merged in this fusion process, resulting in feature redundancy. If maps are decoded on this redundant feature, truly reliable instance features will be lost in each frame. Therefore, we consider that temporal fusion information is beneficial for dynamic target detection, but it lacks the ability to merge static map elements.
@@ III-B1 Vector Point PreSelection Module
$coord_{xy}=Norm(TopK(H_{w}(Map_{inst}))),$ (3)
@@
$pos_{insfea}=S_{a}(F_{pos},coord_{xy}).$ (4)
@@
Finally, the preselected vector points $coord_{xy}$ and the feature $pos_{fea}{=}pos_{insfea}{+}pos_{geofea}$ are used as the initial reference point $P_{m}$ and the point query $Q_{P_{m}}$ to decode each map instance. Note that the query $Q{=}Q_{m}{+}Q_{P_{m}}$ is sent into the deformable DETR where $Q_{m}$ is selected as a random initial variable.
@@
2:  def FindNeg (set, label):{Find the number with this label in this set.}
@@ III-B2 Contrastive Learning of Aggregated Instance Feature
Subsequently, an elaborate selection mechanism is served for positive and negative samples. In adjacent frames of temporal series, changes in the dynamic environment may cause differences in features between instances with a similar position. In particular, parts of instance features are sparse due to the occlusion. However, through the learning process of positive samples, sparse feature expression can be enhanced from relatively rich features. Thus, we choose historical instances within a certain range around the ‘$v$’ instance as positive samples. The negative samples are selected from the negative label, allowing the features of instances with the same labels to be pulled closer together. The negative label refers to a category other than the label of ‘$v$’ itself. The number of negative samples with each negative label is fixed. Several of the highest instance scores in the current frame with negative labels are selected as negative samples to improve learning efficiency. Finally, the contrastive loss of each class is defined as follows:
@@ III-C Temporal Map Consistent Learning
Conventionally, random grids in temporal grid maps aligned by the ego-motion have a similar occupancy probability. Under the guidance of this idea, an MO Loss is designed to control the consistency between temporal grid maps. The process of aligning is achieved through the motion transformation, as explained in Eq. 9:
$Tr$ $=[R|T],$ (9)
$V2GMap_{t-i}^{{}^{\prime}}$ $=Tr(V2GMap_{t-i}),$
where $R$ and $T$ are rotation and translation for the ego respectively. Finally, the occupancy loss is defined as Eq. 10:
@@ III-D Training Loss
Direction Loss: Based on the point matching, this considers the connecting lines between adjacent points. Supervision is done by the cosine similarity of the connecting lines $e_{ij}$ and $eg_{ij}$:
@@ IV-C Implementation Details
HDMapNet [10] EB0 MLP - 30 28.3 7.1 32.6 22.7
VectorMapNet [35] R50 IPM - 110 47.3 36.1 39.3 40.9
InstaGram [42] EffiNet-B0 MLP - 30 40.8 30.0 39.2 36.7
MapTR [11] R50 Transformer 1 24 51.5 46.3 53.1 50.3
MapVR [43] R50 Transformer 1 24 54.4 47.7 51.4 51.2
PivotNet [41] R50 Transformer 4 30 56.5 56.2 60.1 57.6
MapTRv2* [37] R50 LSS - 24 61.4 57.8 60.4 59.9
DTCLMapper-b R50 Transformer 1 24 53.9 51.2 55.8 53.6
DTCLMapper-l R50 LSS - 24 62.6 59.7 63.4 61.9
@@ IV-D Comparison with State-of-the-Arts
Results on nuScenes: Table I shows the comparison results on the nuScenes dataset. To conduct a comprehensive comparison, we verify our methods in two different transformation manners, i.e., DTCLMapper-b using a single layer of transformer, and DTCLMapper-l using LSS. Compared to the transformer-based MapTR and MapVR, the mAP ($53.6\%$) of DTCLMapper-b is higher, yielding respective ${+}3.3\%$ and ${+}2.4\%$ gains. While using LSS, our DTCLMapper-l model can outperform LSS-based MapTRv2 and obtain the highest mAP score of $61.9\%$. Compared to other models [10, 35, 42] learning instance features from a single frame, our DTCLMapper obtains significant gains thanks to the proposed methods enhancing the representation of weak instance features through temporal consistency. Apart from quantitative results, we further conduct qualitative analysis. The visualization results are shown in Fig. 5. Compared to MapTR and MapTRv2 methods, our method performs more consistently in different driving scenarios and yields accurate HD maps. Additionally, the grid maps rasterized from the vectorized maps (resolution of $0.15m$) and a merged map fused from multi-frames are also represented to better showcase the generated HD maps. The proposed rasterization method can accurately describe the occupancy form of the vector map. The experiment results prove the advance of our DTCLMapper in constructing HD Map by using the proposed consistency learning methods.
Method Epo $\rm{AP_{d}}$ $\rm{AP_{p}}$ $\rm{AP_{r}}$ mAP
HDMapNet [10] 6 5.7 13.1 37.6 18.8
VectorNetMapNet [35] 24 6.1 38.3 39.2 37.9
MapVR [43] 6 60.0 54.6 58.0 57.5
MapTR* [11] 6 62.7 55.0 58.5 58.8
MapTRv2* [37] 6 68.8 61.3 63.4 64.5
DTCLMapper-b 6 64.7 56.3 60.1 60.4
DTCLMapper-l 6 69.4 61.9 64.1 65.1
@@ IV-E Ablation Study
Aside from comparing against state-of-the-art methods, we further ablate different components of our DTCLMapper framework. Firstly, the core blocks (VPPSM, AIFCL, MCL, and PE) are explored separately. Then, the parameters in consistent learning are analyzed. We also visualize the preselected vector points for comparison. The map occupancy loss is further compared to other loss functions. Besides, other temporal fusion methods and the merging of temporal maps are discussed separately.
TABLE III: Ablation results on proposed modules.
VPPSM AIFCL MCL PE mAP
50.3
51.6
52.5
52.9
53.6
Analysis of Core Blocks: Table III represents the experiment results of the core blocks. The baseline is a model based on transformers [11]. Subsequent experiments are realized on this basis. We start with the experiment with APPSM and AIFCL, as core modules in the ICL component. Each of them gives the model an accuracy boost (${+}1.3\%$ and ${+}0.9\%$). MCL component brings an improvement of ${+}0.4\%$ to the model. Finally, a geometric position encoder is added to VPPSM, which leads to the best result of $53.6\%$ in mAP. In addition, we analyze the effect of the view transformer module, as shown in Table VI. The result indicates that LSS with depth supervised can promote the transformation between perspective and BEV view, demonstrating the guiding effect of depth supervised.
Visualization of PreSelected Vector Points: Fig. 7 presents a visualization of the preselected vector points. We select the reference point from the initial layer, the second layer, and the fourth layer of optimization to realize comparison. It can be found that compared with the random initialization setting, the initial vector points are near instances in the proposed VPPSM. Observing the distribution of the vector points in the subfigures at different layers, compared with another method, the vector points can regress to the correct instance more quickly and accurately under this initialization setting. Moreover, as shown in the subgraph of the predicted vectorized maps, the proposed method can obtain more accurate map instances.
TABLE IV: Results about different methods of the view transformer.
@@
TABLE V: Results about the number of key embedding in IL.
The max number of ‘$v$ $\rm{AP_{d}}$ $\rm{AP_{p}}$ $\rm{AP_{r}}$ mAP
@@
TABLE VI: Ablation results on Occupancy Loss.
@@
Analysis of Map Occupancy Loss: The MO loss constrains the geometric position of all instances by their occupancy state globally, which is a self-supervision without labels. Three loss functions are compared to construct our optimal MO loss, as shown in Table VI. Here, the resolution of a grid map is $0.15m$, which is a default value. It shows that L1-loss is the best choice, which can play the role of the MO loss precisely.
Analysis of Temporal Fusion: The role of temporal information is a core topic in BEV research. In this part, we analyze the influence of mainstream temporal fusion methods on static map detection. Two methods are selected as research objects, i.e., self-attention [8] and GRU [13]. In the experiment, aligned features from consecutive frames are merged in the BEV feature layer first. Note that the first frame without historical information regards this frame information as history. Table VII shows the result that two temporal fusion methods produce a negative effect on detecting map elements. Upon examining the BEV features of different methods (as depicted in Fig. 2), it becomes evident that the features of temporal fusion methods are ambiguous around map instances, consequently leading to reduced efficiency in instance point regression. In contrast, our method exhibits distinct features surrounding instances. This resonates with our findings concerning different methods in Fig. 1. An indiscriminate fusion method leads to redundant features, particularly for static and similar features, thus impacting the accurate learning of essential elements. However, the proposed consistent learning, concentrated on instances, can improve map detection. It can drive the perspective of studying temporal problems from the view of consistency.
TABLE VII: Results about different temporal fusion methods.
@@
TABLE VIII: Results about the efficiency of different methods. “VT” mans the translated methods between perspective and BEV views. “Trans” means Transformer construct. “SA” means self-attention.
@@
FPS 4.3 4.1 3.6 4.1 4.1 3.8
Analysis of Efficiency: Update frequency is important in studying online vectorized map construction. Table VIII presents the analysis of the efficiency of different methods, including temporal fusion methods. The methods based on Transformer are more efficient compared with the LSS methods. Our approach demonstrates comparable efficiency to state-of-the-art methods while achieving top accuracy.

### v2

@@ III-B Temporal Instance Consistent Learning
As mentioned above, most previous works [8, 23, 5], use temporal information to improve the richness of visual features in BEV space. In their works, global BEV features, as the basic unit, are analyzed through convolution or attention mechanism to get the global weight matrix in temporal frames. Then, this matrix is leveraged to fuse the missing features of this frame from the historical BEV features. During the BEV feature self-learning process from scratch, minimal differences in static objects across consecutive frames within short temporal sequences (as analyzed at the top of Fig. 2) hinder the extraction of reliable deep features from perspective images via existing methods. Specifically, subtle differences can guide modules to repetitively learn overlapping features from similar characteristics, potentially causing feature redundancy phenomena. If maps are decoded on this redundant feature, truly reliable instance features will be lost in each frame. Therefore, we consider that temporal fusion information is beneficial for dynamic target detection, but it lacks the ability to merge static map elements.
@@ III-B1 Vector Point PreSelection Module
$W_{map}$ $=H_{w}(Map_{inst}),$ (6)
$coord_{xy}$ $=Norm(TopK(Flatten(W_{map})),$
@@
$coord_{xy}^{{}^{\prime}}$ $=coord_{xy}\cdot[H,W]^{\rm T},$ (7)
$coord_{xy}^{{}^{\prime\prime}}$ $=coord_{xy}^{{}^{\prime}}[0]+coord_{xy}^{{}^{\prime}}[1]*W,$
$pos_{insfea}$ $=S_{a}(Flatten(F_{pos}),coord_{xy}^{{}^{\prime\prime}}).$
@@
Finally, the preselected vector points $coord_{xy}$ and the feature $pos_{fea}{=}pos_{insfea}{+}pos_{geofea}$ are used as the initial reference point $P_{m}$ and the point query $Q_{P_{m}}\in\mathbb{R}^{N{\times}N_{p}{\times}C}$ to decode each map instance. Note that the query $Q{=}Q_{m}{+}Q_{P_{m}}$ is sent into the deformable DETR where $Q_{m}\in\mathbb{R}^{N{\times}C}$ is initialed as a random variable and extended to be consistent with $Q_{P_{m}}$ by repeat operation.
@@
2:  def FindNeg (set, label):{Find the number without this label in this set.}
@@ III-B2 Contrastive Learning of Aggregated Instance Feature
Subsequently, an elaborate selection mechanism is served for positive and negative samples. In adjacent frames of temporal series, changes in the dynamic environment may cause differences in features between instances with a similar position. In particular, parts of instance features are sparse due to the occlusion. However, through the learning process of positive samples, sparse feature expression can be enhanced from relatively rich features. Thus, we choose historical instances $Ins_{his}$ within a certain range around the ‘$v$’ instance as positive samples $Sam_{pos}$. The negative samples $Sam_{neg}$ are selected from the negative label, allowing the features of instances with the same labels to be pulled closer together. The negative label refers to a category other than the label of ‘$v$’ itself. The number of negative samples with each negative label is fixed. Several of the highest instance scores in the current frame with negative labels are selected as negative samples to improve learning efficiency.
$\left\{\begin{array}[]{ll}L_{p},&label(Ins_{his})=label(v)\\ L_{n},&label(Ins_{his})!=label(v),\end{array}\right.\vspace{-0.1em}$ (9)
$Sam_{pos}=min_{i}(Sqrt(X_{L_{p_{i}}}-X_{v})^{2}+(Y_{L_{p_{i}}}-Y_{v})^{2})),\vspace{-0.5em}$ (10)
$Sam_{neg}=\mathop{max}\limits_{k}(L_{0},\ldots L_{n}),\vspace{-0.3em}$ (11)
where $L_{p}$ and $L_{n}$ are chosen by the label. $k$ is a fixed setting. Finally, the contrastive loss of each class is defined as follows:
@@ III-C Temporal Map Consistent Learning
Conventionally, random grids in temporal grid maps aligned by the ego motion have a similar occupancy probability. Under the guidance of this idea, an MO Loss is designed to control the consistency between temporal grid maps. Given the coordinates $W_{xy}{=}{[X_{0},Y_{0},0],\ldots,[X_{w},Y_{w},0]}$ in the world coordinate system, the process of aligning is achieved through the motion transformation, as explained in Eq. 15 and Eq. 16.
$XY^{{}^{\prime}}=[R|T]_{3\times 3}\cdot W_{xy},$ (15)
$V2GMap_{t-i}^{{}^{\prime}}=S_{a}(Flatten(V2GMap_{t-i}),XY^{{}^{\prime}}),$ (16)
where $R$ and $T$ are rotation and translation for the ego, respectively. $S_{a}$ is the sampling function, according to $XY^{{}^{\prime}}$ sampling points in the $V2GMap_{t-i}$ to obtain the occupied state, which is consistent with the previous definition. Finally, the occupancy loss is defined as Eq. 17:
@@
HDMapNet [6] EB0 MLP - 30 28.3 7.1 32.6 22.7
VectorMapNet [37] R50 IPM - 110 47.3 36.1 39.3 40.9
InstaGram [44] EffiNet-B0 MLP - 30 40.8 30.0 39.2 36.7
MapTR [7] R50 GKT 1 24 51.5 46.3 53.1 50.3
MapVR [45] R50 GKT 1 24 54.4 47.7 51.4 51.2
PivotNet [43] R50 BEVFormer 4 30 56.5 56.2 60.1 57.6
MapTRv2* [39] R50 LSS - 24 61.4 57.8 60.4 59.9
DTCLMapper-b R50 GKT 1 24 53.9 51.2 55.8 53.6
DTCLMapper-l R50 LSS - 24 62.6 59.7 63.4 61.9
@@ III-D Training Loss
Direction Loss: Based on the point matching, this considers connecting lines between adjacent points. Supervision is done by the cosine similarity of the connecting lines $e_{ij}$ and $eg_{ij}$:
@@ IV Experiment
Method Epo $\rm{AP_{d}}$ $\rm{AP_{p}}$ $\rm{AP_{r}}$ mAP
HDMapNet [6] 6 5.7 13.1 37.6 18.8
VectorMapNet [37] 24 6.1 38.3 39.2 37.9
MapVR [45] 6 60.0 54.6 58.0 57.5
MapTR* [7] 6 62.7 55.0 58.5 58.8
MapTRv2* [39] 6 68.8 61.3 63.4 64.5
DTCLMapper-b 6 64.7 56.3 60.1 60.4
DTCLMapper-l 6 69.4 61.9 64.1 65.1
@@ IV-D Comparison with State-of-the-Arts
Results on nuScenes: Table I shows the comparison results on the nuScenes dataset. To conduct a comprehensive comparison, we verify our methods in two different transformation manners, i.e., DTCLMapper-b using a single layer of transformer, and DTCLMapper-l using LSS. Compared to the transformer-based MapTR and MapVR, the mAP ($53.6\%$) of DTCLMapper-b is higher, yielding respective ${+}3.3\%$ and ${+}2.4\%$ gains. While using LSS, our DTCLMapper-l model can outperform LSS-based MapTRv2 and obtain the highest mAP score of $61.9\%$. Compared to other models [6, 37, 44] learning instance features from a single frame, our DTCLMapper obtains significant gains thanks to the proposed methods enhancing the representation of weak instance features through temporal consistency. Apart from quantitative results, we further conduct qualitative analysis. The visualization results are shown in Fig. 5. Compared to MapTR and MapTRv2 methods, our method performs more consistently in different driving scenarios and yields accurate HD maps. Additionally, the grid maps rasterized from the vectorized maps (resolution of $0.15m$) and a merged map fused from multi-frames are also presented to prove the validity of the rasterization method.
Figure 7: Visualization of long temporal fusion on the nuScenes validation dataset. On the left is the comparison against StreamMapNet [50]. On the right is the comparison against MapTracker [51].
TABLE III: Results on the different memory buffer strategies. * represents that the data is realized at the same implementation by open source.
Method Epo training bsz mAP FPS
StreamMapNet* [50] 30 8 60.3 16.3
DTCLMapper-Stream 30 8 61.4 15.9
MapTracker* [51] 30 (4,16,8) 65.4 12.8
DTCLMapper-Tracker 30 (4,16,8) 66.9 12.3
TABLE IV: Ablation results on proposed modules.
VPPSM AIFCL MCL PE mAP FPS
50.3 16.5
51.6 16.1
52.5 16.1
52.9 16.1
53.6 16.0
@@
Results of Different Memory Buffer Strategies: Table III shows the results of different memory buffer strategies on the nuScenes dataset. Two methods are chosen, including StreamMapNet [50] and MapTracker[51]. The proposed temporal consistency learning contributes to improving the mapping quality of these two existing temporal memory buffer strategies, achieving scores of $61.0\%$ and $66.9\%$ in mAP, respectively. Moreover, in terms of efficiency, we remain at a comparable level with these methods. Fig. 7 further demonstrates the beneficial effects of our proposed method on map generation, where our approach accurately identifies instance elements and produces a complete HD map.
@@ IV-E Ablation Study
Aside from comparing against state-of-the-art methods, we further ablate different components of our DTCLMapper. Firstly, the core blocks (VPPSM, AIFCL, MCL, and PE) are explored separately. Then, the parameters in consistent learning are analyzed. Then, the basic framework and matching loss are analyzed. We also visualize the preselected vector points for comparison. The map occupancy loss is further compared to other loss functions. Besides, other temporal fusion methods and the merging of temporal maps are discussed separately.
TABLE V: Results about different methods of the view transformer.
@@
Figure 9: Visual analysis of MO Loss. “With” means MO Loss is used, whereas “Without” means it is not used. Two comparison experiments in different scenarios are shown. Rasterized maps are derived through vectorized instances, while residual maps represent the difference between maps from consecutive frames.
Analysis of Core Blocks: Table IV represents the experiment results of the core blocks. The baseline is a model based on transformers [7]. Subsequent experiments are realized on this basis. We start with the experiment with APPSM and AIFCL, as core modules in the ICL component. Each of them gives the model an accuracy boost (${+}1.3\%$ and ${+}0.9\%$). MCL component brings an improvement of ${+}0.4\%$ to the model. Finally, a geometric position encoder is added to VPPSM, which leads to the best result of $53.6\%$ in mAP. Meanwhile, we present the inference speed of each module. The proposed method only increases computation in the VPPSM and PE modules. From the results, our method shows comparable efficiency to the baseline. In addition, we analyze the effect of the view transformer module. In contrast to transformer-based for 3D-to-2D BEV generation, LSS leverages perspective depth estimation in conjunction with intrinsic and extrinsic parameter projections to derive BEV features. Consequently, depth estimation plays a pivotal role in BEV generation within the LSS approach. As shown in Table V, LSS without depth supervision performs less effectively than the transformer-based method, but LSS with depth supervision can promote the transformation between perspective and BEV view, demonstrating the guiding effect of depth supervision.
TABLE VI: The ablation experiments on framework.
Epo 12 12 12 12 12
Layer 4 6 7 6 6
Inst 50 50 50 40 30
mAP 49.0 53.1 53.2 48.3 52.1
TABLE VII: The ablation experiments on hybrid loss.
Epo 12 12 12 12 12
T 120 240 300 250 200
K 6 6 6 5 4
mAP 49.5 51.5 53.1 52.4 52.0
TABLE VIII: Results about the number of key embedding in IL.
The max number of ‘v’ $\rm{AP_{d}}$ $\rm{AP_{p}}$ $\rm{AP_{r}}$ mAP
@@
Analysis of the Number of Decoder Layer: In Table VIII, we tested the impact of different numbers of decoder layers on accuracy. While a $7$-layer decoder achieves the highest accuracy, it only improves by $0.1\%$ compared to a $6$-layer decoder and introduces a significant increase in parameter count. Therefore, we opt for a $6$-layer decoder structure.
Analysis of the Number of Instance Query: Under the constrained experimental conditions, we tested instance numbers of $30$ per frame, $40$ per frame, and $50$ per frame. Table VIII presents the experimental results, showing that the setting of $50$ instances per frame yields the highest gain.
Analysis of One-to-Many Loss: We also conducted ablation experiments on two parameters, $K$ (indicating the number of query repetitions) and $T$ (representing the additional number of queries during training). In Table VIII, it can be observed that $K{=}6$ and $T{=}300$ reach the best accuracy.
Visualization of PreSelected Vector Points: Fig. 8 presents a visualization of the preselected vector points. We select the reference point from the initial layer, the second layer, and the fourth layer of optimization to realize comparison. It can be found that compared with the random initialization setting, the initial vector points are near instances in the proposed VPPSM. Observing the distribution of the vector points in the subfigures at different layers, compared with another method, the vector points can regress to the correct instance more quickly and accurately under this initialization setting. Moreover, as shown in the subfigures of the predicted vectorized maps, the proposed method can obtain more accurate map instances.
TABLE IX: Ablation results on Occupancy Loss.
@@
TABLE X: Results about different temporal fusion methods.
@@
TABLE XI: Results about the efficiency of different methods. “VT” mans the translated methods between perspective and BEV views. “Trans” means Transformer construct. “SA” means self-attention.
@@
FPS 16.5 16.4 14.7 15.7 16.0 15.3
Analysis of Map Occupancy Loss: The MO loss constrains the geometric position of all instances by their occupancy state globally, which is a self-supervision without labels. Three loss functions are compared to construct our optimal MO loss, as shown in Table IX. Here, the resolution of a grid map is $0.15m$, which is a default value. It shows that L1-loss is the best choice, which can play the role of the MO loss precisely. Fig. 9 further illustrates the visual analysis with or without using MO Loss, and the results further indicate the proposed MO Loss effectively narrows the gap between consecutive frames and produces higher-quality maps.
Analysis of Temporal Fusion: The role of temporal information is a core topic in BEV research. In this part, we analyze the influence of mainstream temporal fusion methods on static map detection. Two methods are selected as research objects, i.e., self-attention [8] and GRU [10]. In the experiment, aligned features from consecutive frames are merged in the BEV feature layer first. Note that the first frame without historical information regards this frame information as history. Table X shows the result that two temporal fusion methods produce a negative effect on detecting map elements. Upon examining the BEV features of different methods (as depicted in Fig. 2), it becomes evident that the features of temporal fusion methods are ambiguous around map instances, consequently leading to reduced efficiency in instance point regression. In contrast, our method exhibits distinct features surrounding instances. This resonates with our findings concerning different methods in Fig. 1. An indiscriminate fusion method leads to redundant features, particularly for static and similar features, thus impacting the accurate learning of essential elements. However, the proposed consistent learning, concentrated on instances, can improve map detection. It can drive the perspective of studying temporal problems from the view of consistency.
Analysis of Efficiency: The update frequency is important in studying online vectorized map construction for autonomous driving. Table XI presents the analysis of the efficiency of different methods which are based on MapTR or MapTRv2. The transformer-based methods are more efficient compared with the LSS methods. Additionally, our approach demonstrates comparable efficiency to state-of-the-art methods while achieving top accuracy.